OpenStudy (anonymous):

What is the prime factorization of 126

terenzreignz (terenzreignz):

Were you taught any method for this?

OpenStudy (anonymous):

Answer : 2 • 3^2 • 7

OpenStudy (anonymous):

no I was not

terenzreignz (terenzreignz):

Okay, then, let's just exhaust it ^^ Can you think of a number that you can divide 126 by? Any number...

OpenStudy (anonymous):

12

terenzreignz (terenzreignz):

No... 126/12 = 10 with a remainder of 6... There mustn't be any remainders... also, stick to using primes...

OpenStudy (anonymous):

2

terenzreignz (terenzreignz):

Okay, good ^_^ What's 126 divided by 2?

OpenStudy (anonymous):

63

terenzreignz (terenzreignz):

Okay. So let's keep that 2. Now... any number that 63 can be divided by?

terenzreignz (terenzreignz):

\[\Large 126 = 2\times 63\] 63 can be divided by...?

OpenStudy (anonymous):

it can not be dvided by 2 right

terenzreignz (terenzreignz):

Nope... it's not even...however, there ARE numbers you can divide it by (with no remainder, of course)

OpenStudy (anonymous):

3

terenzreignz (terenzreignz):

Okay, great, and what's 63 divided by 3?

OpenStudy (anonymous):

21

terenzreignz (terenzreignz):

Okay, let's recap then. \[\Large 126 = 2\times \color{red}{63}= 2\times \color{red}{3\times 21}\] So far so good?

OpenStudy (anonymous):

okay

OpenStudy (loser66):

yes, sir

terenzreignz (terenzreignz):

What about 21, then? Can THAT be divided by anything?

OpenStudy (anonymous):

3

terenzreignz (terenzreignz):

That's right, and 21 divided by 3 is...?

OpenStudy (anonymous):

7

terenzreignz (terenzreignz):

Okay, so we get.. \[\Large 126 = 2\times 63 =2 \times 3 \times \color{green}{21}= 2 \times 3 \times \color{green}{3 \times 7}\] Finally we have 7... can THAT be divided by anything?

OpenStudy (anonymous):

no

terenzreignz (terenzreignz):

Of course not, because 7 is prime ^_^ So, all these factors are prime, now, so THIS \[\Large 126 = \color{blue}{2\times 3\times3 \times 7}\] IS the prime factorisation of 126. Get it? Got it? Good. ^_^

OpenStudy (anonymous):

yes thank you